### v4

# Robust Deterministic Policy Gradient for Disturbance Attenuation and Its Application to Quadrotor Control

Abstract
Practical control systems pose significant challenges in identifying optimal control policies due to uncertainties in the system model and external disturbances. While $H_{\infty}$ control techniques are commonly used to design robust controllers that mitigate the effects of disturbances, these methods often require complex and computationally intensive calculations. To address this issue, this paper proposes a reinforcement learning algorithm called Robust Deterministic Policy Gradient (RDPG), which formulates the $H_{\infty}$ control problem as a two-player zero-sum dynamic game. In this formulation, one player (the user) aims to minimize the cost, while the other player (the adversary) seeks to maximize it. We then employ deterministic policy gradient (DPG) and its deep reinforcement learning counterpart to train a robust control policy with effective disturbance attenuation. In particular, for practical implementation, we introduce an algorithm called robust deep deterministic policy gradient (RDDPG), which employs a deep neural network architecture and integrates techniques from the twin-delayed deep deterministic policy gradient (TD3) to enhance stability and learning efficiency. To evaluate the proposed algorithm, we implement it on an unmanned aerial vehicle (UAV) tasked with following a predefined path in a disturbance-prone environment. The experimental results demonstrate that the proposed method outperforms other control approaches in terms of robustness against disturbances, enabling precise real-time tracking of moving targets even under severe disturbance conditions.
keywords:
Deep reinforcement learning, Two players zero-sum game, Unmanned aerial vehicles
1 Introduction
In recent years, deep reinforcement learning (DRL) has shown remarkable success in finding optimal policies across various domains, including games [1, 2] and control tasks [3, 4]. In particular, various model-free DRL algorithms, such as, deep deterministic policy gradient (DDPG) [3], twin delayed DDPG (TD3) [5], soft actor-critic (SAC) [6], and proximal policy optimization (PPO) [7], have demonstrated the ability to generate effective control inputs without requiring explicit system models. However, a critical challenge arises from the discrepancy between the simulation environments used for training and the real-world environment. Therefore, developing DRL policies that are robust to unseen disturbances is crucial for ensuring reliability in real-world applications. To address this challenge, in this paper, we introduce adversarial learning, where the agent is trained in the presence of disturbances. By exposing the policy to worst-case scenarios during training, the agent can improve its robustness and generalization to unseen environments.
In this paper, we present the so-called robust deterministic policy gradeint (RDPG) algorithm which enhances the robustness of DPG algorithm from the concept of the $H_{\infty}$ control and the two-player zero-sum dynamic game perspectives. The $H_{\infty}$ control method is a robust control technique that aims to design an optimal controller that minimizes the impact of disturbances on system performance. A well-established result in control theory shows that, under certain assumptions, the $H_{\infty}$ control problem can be reformulated as a two-player zero-sum dynamic game, or equivalently, a min-max optimization problem. In this game-theoretic formulation, the controller (first player) seeks to minimize a given cost function, while the adversary (second player) simultaneously tries to maximize it by injecting disturbances [8].
Building upon this theoretical foundation, we transform the $H_{\infty}$ control problem into a two-player zero-sum dynamic game between a user agent (controller) and an adversarial agent (disturbance generator). We then propose a reinforcement learning approach based on the deterministic policy gradient (DPG) method to simultaneously train both agents. Each agent learns its own policy: the user aims to minimize the objective function, while the adversary aims to maximize it. As a result of this zero-sum game framework, the user learns a control policy that is inherently robust against worst-case disturbances. Furthermore, we extend this idea to high-dimensional continuous control settings by incorporating it into a deep reinforcement learning framework. The resulting algorithm, named robust deep deterministic policy gradient (RDDPG), combines the core idea of RDPG with TD3.
Additionally, we aim to develop a robust control strategy for quadrotors, a widely used unmanned aerial vehicle (UAV) in various applications. The ability to maintain robust control of quadrotors is particularly important in dynamic and uncertain environments, where tasks such as tracking moving targets or following predefined waypoints are required. We demonstrate that the proposed RDDPG method effectively determines optimal control inputs in the presence of external disturbances. Furthermore, it outperforms existing DRL-based control approaches by achieving lower cost under severe disturbance conditions.
In conclusion, the main contributions can be summarized as follows:
  1. 1.
    We propose the robust deterministic policy gradient (RDPG) algorithm, which enhances robustness against external disturbances by integrating concepts from $H_{\infty}$ control and two-player zero-sum dynamic games into the actor-critic framework.
  2. 2.
    We develop a novel deep reinforcement learning algorithm, named RDDPG, which integrates RDPG into the TD3 framework to enable scalable and robust learning in continuous control tasks.
  3. 3.
    We apply the proposed RDDPG algorithm to the robust tracking control of quadrotors and conduct comprehensive numerical simulations. The results demonstrate that RDDPG significantly outperforms several state-of-the-art DRL methods, including DDPG, SAC, and TD3.
2 Related works
The $H_{\infty}$ control methods are effective in designing robust controllers against disturbances and have demonstrated satisfactory performance in nonlinear dynamic systems [9, 10, 11]. However, these methods require precise model linearization and incur high computational costs due to the need to solve Hamilton-Jacobi-Isaacs (HJI) equations. To address these limitations, alternative approaches leverage deep reinforcement learning (DRL) to determine optimal control policies without the need for solving nonlinear HJI equations.
Recent efforts have extended the ideas of $H_{\infty}$ control and two-player zero-sum games into the DRL framework to enhance robustness under various scenarios. For instance, [12] developed a robust DRL approach to bridge the gap between different environments for quadcopter control tasks. Their method primarily addresses model uncertainties rather than external disturbances by adopting an action-robust reinforcement learning framework [13], where an adversary injects noise directly into the agent’s actions to minimize the agent’s reward. In contrast, our method explicitly introduces external disturbances into the environment dynamics and formulates the learning process as a two-player zero-sum game between the agent and the adversary. By doing so, our approach enables the agent to learn a control policy that is robust to worst-case disturbance scenarios, rather than merely compensating for action-level noise. In [14], the authors proposed a novel approach that models robust locomotion learning as an interaction between the locomotion policy and a learnable disturbance generator. The disturbance is conditioned on the robot’s state and produces appropriate external forces to facilitate policy learning. To address modeling errors and discrepancies between training and testing conditions, robust adversarial reinforcement learning (RARL) [15] was proposed. RARL trains the agent to perform effectively in the presence of a destabilizing adversary that applies disturbance forces to the system. The agent and the adversary are concurrently trained using the trust region policy optimization (TRPO) algorithm, with the agent learning to maximize the expected reward while the adversary learns to minimize it. Although the underlying concept is similar, their approach differs significantly from ours. Both methods adopt stochastic policy frameworks, such as PPO and TRPO, and they do not impose explicit constraints on the adversary, which may lead to the generation of excessively strong or unrealistic disturbances during training. In contrast, our method is based on deterministic policy gradients and explicitly incorporates the $H_{\infty}$ control perspective into a two-player zero-sum game formulation, enabling more stable and practical learning in continuous control tasks.
3 Preliminaries
3.1 $H_{\infty}$ Control
Let us consider the discrete time nonlinear discrete-time system
$x_{k+1}$ $=f(x_{k},u_{k},w_{k},v_{k})$ (1)
$y_{k}$ $=g(x_{k},u_{k},w_{k})$
where $x_{k}\in\mathbb{R}^{n}$ is the state, $u_{k}\in\mathbb{R}^{m}$ is the control input, $w_{k}\in\mathbb{R}^{p}$ is the disturbance, and $v_{k}\in\mathbb{R}^{q}$ is the process noise at time step $k$. Using the state-feedback controller $u=\pi(x)$, the system can be reduced to the autonomous closed-loop system
$x_{k+1}$ $=f(x_{k},\pi(x_{k}),w_{k},v_{k})$
$y_{k}$ $=g(x_{k},\pi(x_{k}),w_{k})$
Assume that the initial state $x_{0}$ is determined by $x_{0}\sim\rho(\cdot)$, where $\rho$ is the initial state distribution. Defining the stochastic processes ${\bf w}_{0:\infty}:=(w_{0},w_{1},\ldots)$ and ${\bf y}_{0:\infty}:=(y_{0},y_{1},\ldots)$, the system can be seen as a stochastic mapping from ${\bf w}_{0:\infty}$ to ${\bf y}_{0:\infty}$ as follows: ${\bf y}_{0:\infty}\sim T_{\pi}(\cdot|{\bf w}_{0:\infty})$, where $T_{\pi}$ is the conditional probability of ${\bf y}_{0:\infty}$ given ${\bf w}_{0:\infty}$. Moreover, defining the $L^{2}$ norm for the general stochastic process ${\bf z}_{0:\infty}:=(z_{0},z_{1},\ldots)$ by
[…]
The $H_{\infty}$ norm of the autonomous system is defined as
[…]
The goal of $H_{\infty}$ control is to design a control policy $\pi$ that minimizes the $H_{\infty}$ norm of the system $T_{\pi}$. However, minimizing $T_{\pi}||_{\infty}$ directly is often difficult in practical implementations. Instead, we typically aim to find an acceptable $\eta>0$ and a controller satisfying $T_{\pi}||_{\infty}\leq\eta$, which is called suboptimal $H_{\infty}$ control problem. Then, the $H_{\infty}$ control problem can be approximated to the problem of finding a controller $\pi$ that satisfies the constraint
[…]
Defining
${J^{\pi}}:=\mathop{\sup}\limits_{{\bf w}_{0:\infty}\neq 0}{\mathbb{E}}\left[{\left.{\sum\limits_{k=0}^{\infty}{(\|y_{k}\|_{2}^{2}-\eta^{2}\|w_{k}\|_{2}^{2})}}\right|\pi}\right]$
the problem can be equivalently written by finding a controller $\pi$ satisfying ${J^{\pi}}\leq 0$. The problem can be solved by
${\pi^{*}}:=\arg\mathop{\min}\limits_{\pi}{J^{\pi}}.$
3.2 Two-player zero-sum dynamic game perspective
According to [8], the suboptimal $H_{\infty}$ control problem can be equivalently viewed as solving a zero-sum dynamic game under certain assumptions. In particular, we consider two decision making agents called the user and the adversary, respectively. They sequentially take control actions $u_{k}$ (user) and $w_{k}$ (adversary) to minimize and maximize cumulative discounted costs, respectively, of the form
${\mathbb{E}}\left[{\sum_{k=0}^{\infty}{\gamma^{k}c(x_{k},u_{k},w_{k})}}\right]$
where $\gamma\in(0,1]$ is the discount factor, $u_{k}$ is the control input of the user, and $w_{k}$ is the control input of the adversary. The two agents have opposing objectives. The user’s primary objective entails minimizing the costs, while the adversary strives to hinder the user’s progress by maximizing the costs. There exist two categories of the two-player dynamic games, the alternating two-player dynamic game and simultaneous two-player dynamic game. In the alternating two-player dynamic game, two agents engaged in decision making take turns in selecting actions to maximize and minimize the cost, respectively. On the other hand, in the simultaneous two-player dynamic game, the two agents take actions simultaneously to maximize and minimize the cost. In this paper, we mainly focus on the simultaneous two-player dynamic game.
We consider the state-feedback deterministic control policy for each agent, defined as $\pi:{\mathbb{R}}^{n}\to{\mathbb{R}}^{m}$ for the user and $\mu:{\mathbb{R}}^{n}\to{\mathbb{R}}^{p}$ for the adversary
$u=\pi(x)\in{\mathbb{R}}^{m},\quad w=\mu(x)\in{\mathbb{R}}^{p},\quad x\in{\mathbb{R}}^{n}$
The corresponding cost function is defined as
$J^{\pi,\mu}:={\mathbb{E}}\left[\left.\sum\limits_{k=0}^{\infty}\gamma^{k}c(x_{k},u_{k},w_{k})\right|\pi,\mu\right]$ (4)
The goal of the dynamic game is to find a saddle-point pair of equilibrium policies $(\pi,\mu)$ (if exists) for which
${J^{{\pi^{*}},\mu}}\leq{J^{\pi*,\mu*}}\leq{J^{\pi,\mu*}},\quad\forall\pi\in\Pi,\quad\forall\mu\in{\rm M}$
where $\Pi$ denotes the set of all admissible state-feedback policies for the user and $\rm M$} denotes the set of all admissible state-feedback policies for the adversary. The above relation equivalently means
${\mu^{*}}={\max_{\mu\in{\rm M}}}{J^{{\pi^{*}},\mu}},\quad{\pi^{*}}={\min_{\pi\in\Pi}}{J^{\pi,{\mu^{*}}}}$
By the min-max theorem, this also implies that
${\max_{\mu\in{\rm M}}}{\min_{\pi\in\Pi}}{J^{\pi,\mu}}={\min_{\pi\in\Pi}}{\max_{\mu\in{\rm M}}}{J^{\pi,\mu}}$
The optimal cost is now defined as
${J^{*}}:={J^{{\pi^{*}},{\mu^{*}}}}$
To show a connection between the dynamic game framework and the $H_{\infty}$ control problem, let us consider the specific per-step cost function
$c(x,u,w)=\left\|{g(x,u,w)}\right\|_{2}^{2}-{\eta^{2}}\left\|w\right\|_{2}^{2}.$
Then the corresponding cost function is then written when $\gamma=1$ as
${J^{\pi,\mu}}={\mathbb{E}}\left[{\left.{\sum\limits_{k=0}^{\infty}{(\|y_{k}\|_{2}^{2}-\eta^{2}\|w_{k}\|_{2}^{2})}}\right|\pi,\mu}\right]$
The user’s optimal policy is
${\pi^{*}}:=\arg{\min_{\pi\in\Pi}}{\max_{\mu\in{\rm M}}}{J^{\pi,\mu}}$
which is structurally very similar to the policy with the $H_{\infty}$ performance in the previous subsection. It is known that under some special conditions such as the linearity, the $H_{\infty}$ suboptimal control policy and the saddle-point policy are equivalent when $J^{*}\leq 0$.
3.3 Bellman equations
Let us define the value function
${V^{\pi,\mu}}(x):={\mathbb{E}}\left[{\left.{\sum\limits_{k=t}^{\infty}\gamma^{k}c(x_{k},u_{k},w_{k})}\right|\pi,\mu,x_{t}=x}\right]$
so that the corresponding cost function is $J^{\pi,\mu}={\mathbb{E}}[V^{\pi,\mu}(x)|x\sim\rho(\cdot)]$. We can also prove that the value function satisfies the Bellman equation
$V^{\pi,\mu}(x)=c(x,\pi(x),\mu(x))+\gamma{\mathbb{E}}[{V^{\pi,\mu}}(x^{\prime})|\pi,\mu]$
where $x^{\prime}$ implies the next state given the current state $x$ and the action taken by $(\pi,\mu)$. Let us define the optimal value function as $V^{*}:=V^{\pi^{*},\mu^{*}}$. Then, the corresponding optimal Bellman equation can be obtained by replacing $(\pi,\mu)$ by $(\pi^{*},\mu^{*})$. Similarly, let us define the so-called Q-function by
${Q^{\pi,\mu}}(x,u,w):={\mathbb{E}}\left[{\left.{\sum\limits_{k=t}^{\infty}{\gamma^{k}c({x_{k}},{u_{k}},{w_{k}})}}\right|{x_{t}}=x,{u_{t}}=u,{w_{t}}=w,\pi,\mu}\right]$
which satisfies the Q-Bellman equation
${Q^{\pi,\mu}}(x,u,w)=c(x,u,w)+\gamma{\mathbb{E}}[{Q^{\pi,\mu}}(x^{\prime},u^{\prime},w^{\prime})|\pi,\mu]$
where $x^{\prime}$ implies the next state given the current state $x$ and action pair $(u,w)$, $u^{\prime}$ means the next action of the user given $x^{\prime}$ and under $\pi$, and $w^{\prime}$ implies the next action of the adversary given $x^{\prime}$ and under $\mu$. Defining the optimal Q-function $Q^{*}:=Q^{\pi^{*},\mu^{*}}$, one can easily prove the optimal Q-Bellman equation
${Q^{*}}(x,u,w)=c(x,u,w)+\gamma{\mathbb{E}}[{Q^{*}}(x^{\prime},u^{\prime},w^{\prime})|{\pi^{*}},{\mu^{*}}]$
4 Proposed method
In this section, we will present the proposed method, robust deterministic policy gradient (RDPG) and its deep reinforcement learning version, robust deep deterministic policy gradient (RDDPG).
Figure  1 illustrates the architecture of the proposed method, RDDPG, based on a two-player zero-sum dynamic game. In this method, two policy, user policy $\pi$ and adversary policy $\mu$, interact within the environment. The user policy generates the action $u_{t}$, while the adversary policy generates the disturbance $w_{t}$, both based on the current state $x_{t}$. These inputs are fed into the environment, which returns the next state $x_{t+1}$ and the cost $c_{t}$ based on the dynamics of the environment. The critic is updated based on $\{x_{t},u_{t},w_{t},x_{t+1},c_{t}\}$ and evaluates the joint performance of the user and adversary. The policy of user and adversary are updated by deterministic policy gradient method to minimize and maximize the objective function $J^{\pi,\mu}$. The detailed procedure is described in the following sections.
Figure 1: Overview of RDDPG. Based on the two-player zero-sum dynamic game, the user policy $\pi$ seeks to minimize the objective function $J^{\pi,\mu}$, while the adversary policy $\mu$ attempts to maximize it by injecting disturbances into the environment.
4.1 Robust deterministic policy gradient
To apply DPG to the two-player zero-sum dynamic games, we consider the parameterized deterministic control policies
$u=\pi_{\theta}(x)\in{\mathbb{R}}^{m},\quad w=\mu_{\phi}(x)\in{\mathbb{R}}^{p},\quad x\in{\mathbb{R}^{n}}$
where $\pi_{\theta}:\mathbb{R}^{n}\rightarrow\mathbb{R}^{m}$ denotes the user’s control policy parameterized by $\theta$ and $\mu_{\phi}:\mathbb{R}^{n}\rightarrow\mathbb{R}^{p}$ denotes the adversary’s control policy parameterized by $\phi$.. The saddle-point problem is then converted to
${\theta^{*}}:=\arg{\min_{\theta}}{\max_{\phi}}{J^{{\pi_{\theta}},{\mu_{\phi}}}}$
which can be solved using the primal-dual iteration
${\theta_{k+1}}=$ ${\theta_{k}}-{\alpha_{\theta}}{\left.{{\nabla_{\theta}}{J^{{\pi_{\theta}},{\mu_{{\phi_{k}}}}}}}\right|_{\theta={\theta_{k}}}},$
${\phi_{k+1}}=$ ${\phi_{k}}+\alpha_{\phi}{\left.{{\nabla_{\phi}}{J^{{\pi_{{\theta_{k+1}}}},{\mu_{\phi}}}}}\right|_{\phi={\phi_{k}}}}$
where $\alpha_{\theta}>0$ and $\alpha_{\phi}>0$ are step-sizes. According to[16], the deterministic gradients can be obtained using the following theorem.
Theorem 1.
The deterministic policy gradients for the user and adversary policies are given by
${\nabla_{\theta}}{J^{{\pi_{\theta}},{\mu_{\phi}}}}={\mathbb{E}}\left[{\left.{{{\left.{{\nabla_{\theta}}{Q^{\pi,{\mu_{\phi}}}}(s,{\pi_{\theta}},{\mu_{\phi}})}\right|}_{\pi={\pi_{\theta}}}}}\right|s\sim\rho}\right]$
and
${\nabla_{\phi}}{J^{{\pi_{\theta}},{\mu_{\phi}}}}={\mathbb{E}}\left[{\left.{{{\left.{{\nabla_{\phi}}{Q^{{\pi_{\theta}},\mu}}(s,{\pi_{\theta}},{\mu_{\phi}})}\right|}_{\mu={\mu_{\phi}}}}}\right|s\sim\rho}\right]$
respectively.
The proof is a simple extension of the DPG theorem in [16], so it is omitted here. A reinforcement learning counterpart to DPG can be obtained by using the samples of the gradients. Next, we will consider the following per-step cost function:
$c(x,u,w)=\tilde{c}(x,u,w)-\eta^{2}\left\|w\right\|_{2}^{2}$
where $\tilde{c}:{\mathbb{R}}^{n}\times{\mathbb{R}}^{m}\times{\mathbb{R}}^{p}\times{\mathbb{R}}^{n}\to{\mathbb{R}}$ is the cost function for the user. Note that the cost function $\tilde{c}$ can be set arbitrarily and is more general than the quadratic cost in the $H_{\infty}$ control problem. This is because the $H_{\infty}$ control problem can be recovered by setting
$\tilde{c}(x,u,w)=\left\|{g(x,u,w,v)}\right\|_{2}^{2}$
In this way, we can consider the two-player zero-sum dynamic game as an extension of the usual reinforcement learning tasks by considering robustness against the adversarial disturbance. The additional disturbance quadratic cost ${\eta^{2}}w^{T}w$ in the cost function discourages the adversary from generating excessively large disturbances, effectively regularizing its behavior and ensuring that the learned policies are robust against realistic worst-case scenarios rather than arbitrary or unbounded inputs.
Let us assume that we found some approximate solutions $\theta^{*}_{\epsilon},\phi^{*}_{\epsilon}$ such that $J^{\pi_{\theta_{\varepsilon}^{*}},\mu_{\phi_{\varepsilon}^{*}}}\leq 0$. Then, this implies that
$\frac{{{\mathbb{E}}\left[{\left.{\sum\limits_{k=0}^{\infty}{\gamma^{k}\tilde{c}({x_{k}},{u_{k}},{w_{k}})}}\right|{\pi_{\theta_{\varepsilon}^{*}}},{\mu_{\phi_{\varepsilon}^{*}}}}\right]}}{{{\mathbb{E}}\left[{\left.{\sum\limits_{k=0}^{\infty}\gamma^{k}\left\|w_{k}\right\|_{2}^{2}}\right|{\pi_{\theta_{\varepsilon}^{*}}},{\mu_{\phi_{\varepsilon}^{*}}}}\right]}}\leq{\eta^{2}}$
4.2 Robust deep deterministic policy gradient
To implement RDPG in high-dimensional continuous control tasks, we incorporate the techniques from twin-delayed deep deterministic policy gradient (TD3)[5], resulting in the robust deep deterministic policy gradient (RDDPG). In particular, we introduce the following networks:
  1. 1.
    The online actor network $\pi_{\theta}$ for the user and the online actor network $\mu_{\phi}$ for the adversary
  2. 2.
    The corresponding target networks $\pi_{\theta^{\prime}}$, $\mu_{\phi^{\prime}}$ for the two online actor network
  3. 3.
    Two online critic networks $Q_{\psi_{1}}(x,u,w),Q_{\psi_{2}}(x,u,w)$
  4. 4.
    The corresponding target networks $Q_{\psi_{1}^{\prime}}(x,u,w)$, $Q_{\psi_{2}^{\prime}}(x,u,w)$
Now, following [5], the actor and critic networks are trained through the following procedures.
4.2.1 Critic update
The critic is trained by the gradient descent step to the loss
$L_{\rm critic}(\psi_{i};B):=$ $\frac{1}{|B|}\sum\limits_{(x,u,w,c,x^{\prime})\in B}{(y-Q_{\psi_{i}}(x,u,w))}^{2},\quad i\in\{1,2\}$
where $B$ is the mini-batch, $|B|$ is the size of the mini-batch, $c$ is the cost incurred at the same time as the state $x$, and $x^{\prime}$ means the next state. Moreover, the target $y$ is defined as
[…]
when $x^{\prime}$ is not the terminal state, and
$y=c$
when $x^{\prime}$ is the terminal state (in the episodic environments), where
$\tilde{u}$ $=\pi_{\theta^{\prime}}(x^{\prime})+{\rm clip}({\epsilon_{1}},-\delta,\delta)\ \ \epsilon_{1}\sim\mathcal{N}(0,\sigma)$
$\tilde{w}$ $=\mu_{\phi^{\prime}}(x^{\prime})+{\rm clip}({\epsilon_{2}},-\delta,\delta)\ ,\ \epsilon_{2}\sim\mathcal{N}(0,\sigma)$
where $\epsilon_{1},\epsilon_{2}$ are random noise vectors added in order to smooth out the target values, $\mathcal{N}(0,\sigma)$ implies the Gaussian distribution with zero mean and variance $\sigma$, and the clip function is added in order to guarantee the boundedness of the control inputs. The critic’s online parameters $\psi_{1},\psi_{2}$ are updated by the gradient descent step to minimize the loss
${\psi_{i}}\leftarrow{\psi_{i}}-{\alpha_{{\rm{critic}}}}{\nabla_{\psi_{i}}}{L_{\rm critic}}({\psi_{i}};B),\quad i\in\{1,2\}$
where $\alpha_{\rm critic}$ is the step-size.
4.2.2 Actor update
The actor networks for the user and adversary are updated using the sampled deterministic policy gradient
$\theta\leftarrow\theta-{\alpha_{\rm actor}}{\nabla_{\theta}}{L_{\rm actor}}(\theta,\phi;B)$
$\phi\leftarrow\phi+{\beta_{\rm actor}}{\nabla_{\phi}}{L_{\rm actor}}(\theta,\phi;B)$
where $\alpha_{\rm actor}$ and $\beta_{\rm actor}$ are the step-sizes, and
$L_{\rm actor}(\theta,\phi;B)=\frac{1}{{|B|}}\sum\limits_{(x,u,w,r,x^{\prime})\in B}{{Q_{{\psi_{1}}}}(x,{\pi_{\theta}}(x),{\mu_{\phi}}(x))}$
Please note the opposite signs of the gradients for the user and adversary updates, which are due to their opposite roles.
4.2.3 Target parameters update
The target parameters are updated through the averaging
$\phi^{\prime}\leftarrow\tau\phi+(1-\tau)\phi^{\prime},\quad\theta^{\prime}\leftarrow\tau\theta+(1-\tau)\theta^{\prime}$
for the actor target parameters and
${\psi_{i}}^{\prime}\leftarrow\tau{\psi_{i}}+(1-\tau){\psi_{i}}^{\prime},\quad i\in\{1,2\}$
for the critic target parameters.
4.2.4 Exploration
For exploration, we can use
$u_{k}=$ $\pi_{\theta}(x_{k})+e_{1}$
$w_{k}=$ $\mu_{\phi}(x_{k})+e_{2}$
where $e_{1},e_{2}\sim\mathcal{N}(0,\sigma)$ are exploration noises [5]. The overall algorithm is summarized in Algorithm 1.
Algorithm 1 RDDPG
1:Initialize the online critic networks $Q_{\psi_{1}},Q_{\psi_{2}}$
2:Initialize the actor networks $\pi_{\theta},\mu_{\phi}$ for the user and adversary, respectively.
3:Initialize the target parameters $\psi_{1}^{\prime}\leftarrow\psi_{1},\psi_{2}^{\prime}\leftarrow\psi_{2}$, $\theta^{\prime}\leftarrow\theta,\phi^{\prime}\leftarrow\phi$
4:Initialize the replay buffer $\mathcal{D}$
[…]
6:  Observe $s_{0}$
[…]
8:   Select actions $u_{k}=\pi_{\theta}(x_{k})+e_{1}$ and $w_{k}=\mu_{\phi}(x_{k})+e_{2}$,
9:   where $e_{1},e_{2}\sim\mathcal{N}(0,\sigma)$ are exploration noises.
10:   Compute the cost $c_{k}:=c(x_{k},u_{k},w_{k})$
11:   Observe the next state $x_{k+1}$
12:   Store the transition tuple $(x_{k},u_{k},w_{k},c_{k},x_{k+1})$ in the replay buffer $D$
13:   Uniformly sample a mini-batch $B$ from the replay buffer $\mathcal{D}$
14:   Update critic network:
${\psi_{i}}\leftarrow{\psi_{i}}-{\alpha_{{\rm{critic}}}}{L_{\rm critic}}({\psi_{i}};B),\quad i\in\{1,2\}$
15:   Update actor networks by the deterministic policy gradient:
$\theta\leftarrow\theta-{\alpha_{\rm actor}}{\nabla_{\theta}}{L_{\rm actor}}(\theta,\phi;B)$
$\phi\leftarrow\phi+{\beta_{\rm actor}}{\nabla_{\phi}}{L_{\rm actor}}(\theta,\phi;B)$
16:   Soft update target networks:
$\theta_{i}^{\prime}\leftarrow\tau\theta_{i}+(1-\tau)\theta_{i}$
$\phi_{i}^{\prime}\leftarrow\tau\theta_{i}+(1-\tau)\phi_{i}$
17:  end for
18:end for
5 Quadrotor application
5.1 Quadrotor dynamics model
The quadrotor dynamics can be described as follows:
$m\mathbf{a}_{k}$ $=m\begin{bmatrix}0\\ 0\\ -g\end{bmatrix}+\mathbf{R}\mathbf{T}_{k}+\mathbf{F}_{D}+w_{k}$ (5)
$\Delta\omega_{k}$ $=I^{-1}(-\omega_{k-1}\times I\omega_{k-1}+\tau_{B})$ (6)
$\mathbf{v}_{k}$ $=\mathbf{v}_{k-1}+T_{s}\mathbf{a}_{k}$ (7)
$\mathbf{p}_{k}$ $=\mathbf{p}_{k-1}+T_{s}\mathbf{v}_{k}$ (8)
$\omega_{k}$ $=\omega_{k-1}+T_{s}\Delta\omega_{k}$ (9)
$\rho_{k}$ $=\rho_{k-1}+T_{s}\omega_{k}$ (10)
where $m$ is the mass of the quadrotor, $g$ is the gravity acceleration, $T_{s}$ is the sampling time, $w_{k}=[w_{k,x},w_{k,y},w_{k,z}]^{T}$ means an external disturbance applied to the quadrotor, $\mathbf{a}_{k}=[a_{k,x},a_{k,y},a_{k,z}]^{T}$ denotes linear acceleration, $\mathbf{v}_{k}=[v_{k,x},v_{k,y},v_{k,z}]^{T}$ represents linear velocity, $\mathbf{p}_{k}=[p_{k,x},p_{k,y},p_{k,z}]^{T}$ describes the position of the quadrotor in the inertial frame, $\Delta\omega_{k}=[\Delta\omega_{k,\phi},\Delta\omega_{k,\theta},\Delta\omega_{k,\psi}]^{T}$ is angular acceleration, $\omega_{k}=[\omega_{k,\phi},\omega_{k,\theta},\omega_{k,\psi}]^{T}$ is the angular velocity, and $\rho_{k}=[\phi_{k},\theta_{k},\psi_{k}]^{T}$ represents the roll, pitch, yaw of the quadrotor in the body frame. Moreover, $\mathbf{R}$ is the rotation matrix from the body frame to the inertial frame, and $\mathbf{T}_{B}$ is the total thrust of the quadrotor along the $z$ axis in the body frame and defined as follows:
$\mathbf{T}_{k}=\sum_{i=1}^{4}T_{k,i}=k_{F}\begin{bmatrix}0\\ 0\\ \sum\text{RPM}_{k,i}^{2}\\ \end{bmatrix}$
where $T_{k,i}$ is the thrust generated by $i$th motor, $k_{F}$ is the thrust coefficient, and $\text{RPM}_{k,i}$ is the speed of $i$th motor at time step $k$. $\mathbf{F}_{D}$ is a drag force, which is generated by motor spinning, defined as
$\mathbf{F}_{D}=-\mathbf{K}_{D}\left(\sum_{i=1}^{4}\text{RPM}_{k,i}\right)\mathbf{v}_{k}$
where $\mathbf{K}_{D}=\text{diag}(K_{D,x},K_{D,y},K_{D,z})$ is a matrix of drag coefficients. Among the four motors, two opposing motors ($1$th and 33rd) rotate clockwise, while the others (22nd and $4$th) rotate counterclockwise. This causes torque $\tau_{k}$ in the roll, pitch, yaw directions
$\tau_{k}=\begin{bmatrix}\tau_{k,\phi}\\ \tau_{k,\theta}\\ \tau_{k,\psi}\end{bmatrix}=\begin{bmatrix}l^{\prime}k_{F}(\text{RPM}_{k,2}^{2}-\text{RPM}_{k,4}^{2})\\ l^{\prime}k_{F}(-\text{RPM}_{k,1}^{2}+\text{RPM}_{k,3}^{2})\\ k_{M}(\text{RPM}_{k,1}^{2}-\text{RPM}_{k,2}^{2}+\text{RPM}_{k,3}^{2}-\text{RPM}_{k,4}^{2})\\ \end{bmatrix}$
where $l^{\prime}=\cos(\frac{\pi}{4})l$ is the scaled arm length, because the motors of the quadrotor are positioned at a 45-degree angle relative to the $x$ and $y$ axes when the yaw angle is 0, $k_{F}$ is the thrust coefficient, and $k_{M}$ is the torque coefficient. The inertia matrix, denoted by the symbol $I$, represents the inertial properties of the quadrotor’s body.
5.2 State space
The state $x_{k}$ of the quadrotor is defined as
$x_{k}=[\mathbf{e}_{k}^{p},\mathbf{e}_{k}^{v},\boldsymbol{\rho}_{k},\boldsymbol{\omega}_{k}]^{T}\in\mathbb{R}^{12}$
where $\mathbf{e}_{k}^{p}$ denotes the position error, $\mathbf{e}_{k}^{v}$ represents the velocity error, $\boldsymbol{\rho}_{k}$ is the Euler angle vector, and $\boldsymbol{\omega}_{k}$ is the angular velocity vector. These quantities are defined as follows:
$\mathbf{e}_{k}^{p}$ $=\mathbf{p}_{k}^{\text{target}}-\mathbf{p}_{k}$
$=[p_{k,x}^{\text{target}}-p_{k,x},\ p_{k,y}^{\text{target}}-p_{k,y},\ p_{k,z}^{\text{target}}-p_{k,z}]^{T}\in\mathbb{R}^{3}$
$\mathbf{e}_{k}^{v}$ $=\mathbf{v}_{k}^{\text{target}}-\mathbf{v}_{k}$
$=[v_{k,x}^{\text{target}}-v_{k,x},\ v_{k,y}^{\text{target}}-v_{k,y},\ v_{k,z}^{\text{target}}-v_{k,z}]^{T}\in\mathbb{R}^{3}$
$\boldsymbol{\rho}_{k}$ $=[\phi_{k},\theta_{k},\psi_{k}]^{T}\in\mathbb{R}^{3}$
$\boldsymbol{\omega}_{k}$ $=[\omega_{k,\phi},\omega_{k,\theta},\omega_{k,\psi}]^{T}\in\mathbb{R}^{3}$
where $\mathbf{p}_{k}^{\text{target}}$ and $\mathbf{p}_{k}$ denote the position vectors of the target and the quadrotor, respectively, and $\mathbf{v}_{k}^{\text{target}}$ and $\mathbf{v}_{k}$ represent their corresponding linear velocity vectors. $\phi_{k}$$\theta_{k}$$\psi_{k}$ denotes roll, pitch, yaw of the quadrotor. As in the method used in DQN[2], we used the last four time steps of the state as input to the neural network to help the network better capture dynamics in the environment.
5.3 Action space
The action of the quadrotor, $u_{k}$, is the speed of four motors $\text{RPM}_{k,i},\ i\in\{1,2,3,4\}$, which is continuous value between 0 and $21713.714$, as follows:
$u_{k}:=[\text{RPM}_{k,1},\text{RPM}_{k,2},\text{RPM}_{k,3},\text{RPM}_{k,4}]^{T}\in\mathbb{R}^{4}.$
The action of the disturbance adversary, $w_{k}$, represents the external forces to the quadrotor along the $x$, $y$, and $z$ axes of the body frame as follows:
$w_{k}:=[w_{k,x},w_{k,y},w_{k,z}]^{T}\in\mathbb{R}^{3}$
5.4 Cost function
The cost function $c(x_{k},u_{k},w_{k})$ is comprised of the cost function for the user $\tilde{c}(x_{k},u_{k})$ and the disturbance quadratic cost $\eta^{2}w_{k}^{T}w_{k}$. The cost function for the user $\tilde{c}(x_{k},u_{k})$ is the sum of the following terms: $\tilde{c}(x_{k},u_{k})=c_{k,p}+c_{k,v}+c_{k,\rho}+c_{l}+c_{tr}$
  1. 1.
    Position error cost $c_{k,p}$ : Position error between the target and quadrotor
[…]
  2. 2.
    Velocity error cost $c_{k,v}$ : Velocity error between the target and quadrotor
[…]
  3. 3.
    Angle cost $c_{k,\rho}$ : Magnitude of roll, pitch, and yaw of the quadrotor
[…]
  4. 4.
    Living cost $c_{l}$ : Constant cost at every step
    $c_{l}=\lambda$
  5. 5.
    Penalty cost $c_{tr}$ : Penalty when the quadrotor fails to track the target
[…]
where the scaling factors $\alpha$, $\beta$, $\epsilon$, and $\zeta$ are positive (i.e., $\alpha,\beta,\epsilon,\zeta>0$), and $\lambda$ is negative (i.e., $\lambda<0$). The specific values used for these parameters are summarized in the Appendix  2. Since $\alpha$ and $\beta$ are positive, the quadrotor agent learns a policy that minimizes position and velocity errors. If the quadrotor’s angles become too large, it becomes more susceptible to disturbances and may flip easily. Therefore, minimizing the angles cost $c_{\rho}$ is crucial for stability. Additionally, the living cost $c_{l}$ is a negative reward that encourages the quadrotor to maintain stability. The penalty cost $c_{tr}$ is a large positive value applied when the quadrotor fails to track the target—specifically, when the distance between the target and the quadrotor exceeds $5$ meters, or when the quadrotor flips with roll or pitch angles exceeding $\frac{\pi}{2}$. The agent learns a policy to avoid receiving this penalty, which can accelerate the initial learning process.
6 Experiments and results
6.1 Experiments setups
For experiments, we used the gym-pybullet-drones environment [17], which is an open-source quadrotor simulator built with Python and the Bullet Physics engine. This environment provided a modular and precise physics implementation, supporting both low-level and high-level control. In the simulator, the Crazyflie 2.0 quadrotor was modeled, and its parameters were summarized in Appendix Table 1.
To demonstrate the robustness of the proposed RDDPG algorithm, we compared it with other model-free reinforcement learning algorithms, including DDPG [3], TD3 [5], and SAC [6]. These methods utilized a single agent that aimed to minimize the cost function $\tilde{c}$. To enhance their robustness against disturbances, random external disturbances were applied during the training process. A detailed summary of the hyperparameters used in training was provided in Appendix 3 and  4.
For evaluation, we generated two trajectories as shown in Fig. 2. During evaluation, we introduced random wind disturbances in the $x$ and $y$ directions. We considered three scenarios where the maximum wind velocity was limited to $\pm 5$, $\pm 10$, and $\pm 15$ [m/s], which were referred to as the $\pm 5$, $\pm 10$, and $\pm 15$ wind speed scenarios in the figures and tables. In each scenario, the initial wind speed was randomly set within a given range and updated every $10$ time steps with normally distributed noise.
Each algorithm was tested over 500 episodes and evaluated based on the mean and standard deviation of the total cost per episode, where the total cost was computed as the sum of the per-step costs $\tilde{c}(x_{k},u_{k})$ over a single episode. In addition, the mean and standard deviation of the position error $\mathbf{e}_{k}^{p}||_{2}^{2}$ between the target and the quadrotor at each timestep were also considered as evaluation metrics.
Traj (1)
Traj (2)
Figure 2: Two evaluation trajectories, denoted as Traj (1) and Traj (2), shown from left to right.
6.2 Results
Figure 3: Mean and standard deviation of the total episode cost. RDDPG achieves the lowest average cost and the smallest variance across all scenarios. This indicates that the RDDPG yields a robust agent that performs consistently even under severe external disturbances.
[…]
By maintaining a consistently low and stable cost across a wide range of disturbance intensities, RDDPG exhibits superior robustness and reliable control performance, even under challenging environmental conditions. This enhanced robustness stems from its adversarial training framework, inspired by $H_{\infty}$ control theory and formulated as a two-player zero-sum game. By optimizing the policy against worst-case disturbances during training, RDDPG learns disturbance-resilient strategies that improve generalization and ensure stable performance under unseen or highly uncertain conditions.
Figure 4: Trajectory results of all algorithms under the $\pm 15$ wind scenario and Traj (1).
Figure 4 illustrates the trajectory tracking performance of each algorithm across varying wind speeds. RDDPG achieves the most accurate trajectory tracking, exhibiting the smallest tracking error compared to the other algorithms. The mean and standard deviation of the position error $\mathbf{e}_{k}^{p}||_{2}^{2}$ between the target and the quadrotor is presented in Figure  5. Across all scenarios, RDDPG consistently achieves lower position errors compared to TD3 and SAC. In the case of DDPG, it exhibits lower position errors under mild wind disturbance scenarios. However, as the disturbance becomes stronger, its performance degrades significantly, resulting in higher errors compared to RDDPG. Furthermore, DDPG shows larger standard deviations, indicating less stable performance across episodes. Detailed numerical results are provided in at Appendix Table  5.
Figure 5: Mean and standard deviation of the position error $\mathbf{e}_{k}^{p}||_{2}^{2}$ between the target and the quadrotor at each time step. RDDPG consistently achieves low error and standard deviation across all scenarios, whereas other methods exhibit unstable performance.
7 Conclusion
This paper proposes robust deterministic policy gradient (RDPG), which combines the concept of the $H_{\infty}$ control problem and a two-player zero-sum dynamic game framework, to overcome the robustness problem of DRL algorithms. In RDPG, the user seeks to minimize the cost function while the adversary attempts to maximize it, thereby enabling the agent to learn a control policy that is robust to external disturbances. Furthermore, we introduce robust deep deterministic policy gradient (RDDPG), which integrates the robustness framework of RDPG with the learning stability of TD3, achieving both improved disturbance resistance and efficient policy learning in continuous control tasks. In order to evaluate the robustness of RDDPG, we implement it on quadrotor tracking tasks under various external disturbances. Experimental results show that RDDPG successfully learns an optimal control policy while maintaining stable performance across different disturbance scenarios. These findings demonstrate that RDDPG outperforms other DRL-based methods in terms of robustness and reliability in real-world-inspired environments.
Appendix A Hyperparameters Used in Training
A.1 Physical parameters of Crazyflies 2.0
Table 1: Physical parameters of the Crazyfile 2.0
Parameter Description Value
m Total mass 0.0270.027[kg]
l Arm length 0.03970.0397[m]
$K_{F}$ Thrust coefficient $3.16\times 10^{-10}$
$K_{M}$ Torque coefficient $7.94\times 10^{-12}$
$I_{xx}$ Principal Moment of Inertia around x axis $1.4\times 10^{-3}$ $\text{kg}\times\text{m}^{2}$]
$I_{yy}$ Principal Moment of Inertia around y axis $1.4\times 10^{-3}$ $\text{kg}\times\text{m}^{2}$]
$I_{zz}$ Principal Moment of Inertia around z axis $2.17\times 10^{-3}$ $\text{kg}\times\text{m}^{2}$]
$K_{D.xy}$ Drag coefficients in $x,y$ axes $9.18\times 10^{-7}$
$K_{D.z}$ Drag coefficients in $z$ axis $10.31\times 10^{-7}$
A.2 Cost function
Table 2: Scaling factors in cost function
Parameter Description Value
$\alpha$ Coefficient for position error $10$
$\beta$ Coefficient for velocity error 11
$\epsilon$ Coefficient for angle error 11
$\lambda$ Living reward 10-10
$\zeta$ Penalty $1000$
A.3 DDPG and TD3
Table 3: Caption
Parameter Description Value
$\alpha_{actor}$ Learning rate for user 0.0001
$\beta_{actor}$ Learning rate for adversary 0.0001
$\alpha_{critic}$ Learning rate for critic 0.001
$|B|$ Batch size 256
$\gamma$ Discounted factor 0.99
$\sigma$ Exploration noise 0.2
A.4 SAC
Table 4: SAC hyper-parameters
Parameter Description Value
$\alpha_{actor}$ Learning rate for user 0.0003
$\alpha_{critic}$ Learning rate for critic 0.003
$|B|$ Batch size 256
$\gamma$ Discounted factor 0.99
$\alpha_{entropy}$ Entropy coefficient 0.99
Appendix B Experiment results
Table 5: The mean and standard deviation of the position error $\mathbf{e}_{k}^{p}||_{2}$ between the target and the quadrotor. Bold indicates the best performance algorithm
$\pm 5$ wind speed $\pm 10$ wind speed $\pm 15$ wind speed
Traj (1) Traj (2) Traj (1) Traj (2) Traj (1) Traj (2)
[…]
Figure 6: In $\pm 15$ m/s wind scenario
Figure 7: In $\pm 15$ m/s wind scenario
Figure 8: In $\pm 15$ m/s wind scenario
Figure 9: In $\pm 15$ m/s wind scenario
Figure 10: In $\pm 15$ m/s wind scenario